# Subjective-Aligned Dateset and Metric for Text-to-Video Quality Assessment

### v2

@@ 4 Subjective-Aligned Text-to-video Metric
Figure 5: Overview framework of T2VQA. Features from text-video alignment and video fidelity perspectives are extracted. After a cross-attention based fusion, an LLM is utilized for regression.
4.1 Text-video Alignment Encoder
@@
Text-video alignment refers to the conformity between the video content and the text description. CLIP [32] and BLIP [27] are models that have a strong ability for zero-shot text-image matching.  [46] first proposes CLIPSim, which uses CLIP to calculate the similarities between text and each frame of the video and then take the average value. However, former works [5, 16, 29, 28] and results in Tab. 2 show that simply using CLIP or BLIP has a low correlation with the authentic subjective scores. Following works of MLLMs such as BLIP-2 [26], InstructBLIP [6], and mPLUG-Owl2 [59], we use the pre-trained BLIP image encoder as the video frame encoder. We freeze the weights of the image encoder that encodes each frame separately. We use the BLIP text encoder as the alignment encoder. The alignment encoder takes the encoded text and video frame as inputs. They interact through the cross-attention module and the encoder eventually outputs a feature representing the text-frame similarity. We concatenate features from each text-frame pair. Given a set of $N$ video frames $\{v_{i}\in\mathbb{R}^{3\times H\times W}\}_{i=1}^{N}$ and the text prompt $t$, we have:
$f_{b}=cat(\{\mathrm{BLIP}(t,v_{i})\}_{i=1}^{N}).$ (2)
4.2 Video Fidelity Encoder
@@
Video fidelity refers to the perception of distortion from spatial and temporal domains. In the spatial domain, common distortion types include blurriness, noises, low/high contrast, etc. Temporal distortions include jitter, stall, motion blur, etc. In this perspective, the task can be seen as a common VQA task. Swin-T [30] has been proven for its excellent ability in various VQA tasks [47, 50]. By using 3D-shifted window-based multi-head self-attention (SW-MSA), Swin-T has a strong ability to analyze videos from spatial and temporal domains. Therefore, we utilize Swin-T as the backbone of the fidelity encoder to extract features that represent video fidelity. Given a video clip $v\in\mathbb{R}^{3\times N\times H\times W}$, we have:
$f_{s}=\mathrm{SWIN}(v).$ (3)
4.3 Feature Fusion
@@
Inspired by BLIP-2 [26] and InstructBLIP [6], after having the features from perspectives of text-video alignment and video fidelity, we design a transformer-based fusion module to fuse those two features. The module includes $N$ blocks with self-attention, cross-attention, and feed-forward layers in each block. The fidelity feature $f_{s}$ first goes through self-attention layers, and then it interacts with the alignment feature $f_{b}$ in cross-attention layers (and every other transformer block). The fusion module helps the model to unify the features from two perspectives to have a more comprehensive understanding of the video characteristics. We initialize the fusion module using $\mathrm{BERT}_{base}$ [19].
4.4 Quality Regression
@@
LLMs have been proven to be competitive on IQA/VQA tasks [52, 53, 64, 54]. Inspired by them, we also utilize an LLM as the quality regression module in T2VQA. We first design a text instruction prompt, i.e., “Please rate the quality of this video.”, to guide the LLM. The encoded instruction and the fused feature are concatenated as the input of the LLM. Following [61, 53], we supervise the LLM to output one among five ITU-standard [17] levels (bad, poor, fair, good, and excellent) to represent the quality of the videos, denoted as <level>. We assign them weights of $1-5$ in order. Since the logit at <level> in LLM is the probability distribution of all tokens, it can be used to represent how the LLM predicts the quality of the video. Therefore, a softmax for each token is calculated and multiplied by its weight. We have the final predicted score as:
$s_{pred}=\sum_{i=1}^{5}i\times softmax(\lambda_{i})=\sum_{i=1}^{5}i\times\frac{e^{\lambda_{i}}}{\sum_{j=1}^{5}e^{\lambda_{j}}},$ (4)
@@
5 Experiments and Results
5.1 Implement Details
5.1.1 Train-test splitting.
@@ 5.1 Implement Details
5.1.2 Training Settings.
We utilize the large model of BLIP for the extraction of text-video alignment feature. We initialize the fidelity encoder using Swin-T pre-trained on Kinetics-400 [18] dataset. For the LLM, we use the 7B model of Vicuna v1.5 [65], which is fine-tuned from Llama 2 [41]. It is worth noting that both BLIP image encoder and LLM are frozen during training.
@@
5.1.3 Loss Function.
Following [47], we use differentiable Pearson Linear Correlation Coefficient (PLCC) and rank loss as loss functions. PLCC is a common criterion used for evaluating the correlation between sequences, while rank loss is introduced to help the model distinguish the relative quality of videos better. The final loss function is defined as:
$L=L_{plcc}+\lambda\cdot L_{rank},$ (5)
@@
5.1.4 Evaluation Metrics.
Along with PLCC, we include Spearman’s Rank-Order Correlation Coefficient (SROCC), Kendall’s Rank-order Correlation Coefficient (KRCC), and Root Mean Square Error (RMSE) as performance criteria. SROCC, PLCC, and KRCC indicate the prediction monotonicity, while RMSE measures the prediction accuracy. Better models should have larger SROCC, KRCC, and PLCC scores, but conversely for RMSE. Before calculating PLCC, we follow the same procedure in [10] to map the objective score to the subject score using a four-parameter logistic function.
Table 2: Performance of the SOTA models and T2VQA on T2VQA-DB. The best-performing model is highlighted in each column. [Keys: S: SROCC, P: PLCC, K: KRCC, R: RMSE, Bold: the best].
Models T2VQA-DB Validation Sora Testing
S $\uparrow$ P $\uparrow$ K $\uparrow$ R $\downarrow$ S $\uparrow$ P $\uparrow$ K $\uparrow$ R $\downarrow$
zero-shot
CLIPSim [32] 0.1047 0.1277 0.0702 21.683 0.2116 0.1538 0.1406 18.316
BLIP [27] 0.1659 0.1860 0.1112 18.373 0.2126 0.1038 0.1515 18.850
ImageReward [58] 0.1875 0.2121 0.1266 18.243 0.0992 0.0415 0.0748 19.494
ViCLIP [45] 0.1162 0.1449 0.0781 21.655 0.2567 0.1844 0.1734 17.982
UMTScore [29] 0.0676 0.0721 0.0453 22.559 0.2594 0.0840 0.1680 19.057
finetuned
SimpleVQA [39] 0.6275 0.6338 0.4466 11.163 0.0340 0.2344 0.0237 16.687
BVQA [23] 0.7390 0.7486 0.5487 15.645 0.4235 0.2489 0.2635 17.164
FAST-VQA [47] 0.7173 0.7295 0.5303 10.595 0.4301 0.2369 0.2939 17.426
DOVER [50] 0.7609 0.7693 0.5704 9.8072 0.4421 0.2689 0.2757 17.182
T2VQA(Ours) 0.7965 0.8066 0.6058 9.0221 0.6485 0.3124 0.4874 16.511
5.2 Performance Comparison
5.2.1 Reference Algorithms.
For performance comparison, we use CLIPSim [46], BLIP [27], ImageReward [58], ViCLIP [45], UMTScore [29], SimpleVQA [39], BVQA [23], FAST-VQA [47], and DOVER [50] as the reference algorithms. DOVER is considered the SOTA VQA method to date. CLIPSim averages the similarity values between the text and each video frame. We adopt the same operation on BLIP and ImageReward to tune IQA metrics into VQA metrics. ViCLIP and UMTScore are metrics designed for measuring text-video alignment. SimpleVQA, BVQA, FAST-VQA, and DOVER are models designed for general VQA tasks. We use the pre-trained weights of CLIPSim, BLIP, and ImageReward for zero-shot testing, and we finetune SimpleVQA, BVQA, FAST-VQA, and DOVER on T2VQA-DB. All results are averaged after ten-fold splitting.
(a) BLIP
(b) SimpleVQA
(c) BVQA
(d) FAST-VQA
(e) DOVER
(f) T2VQA
Figure 6: Scatter plots of the predicted scores vs. MOSs. The curves are obtained by a four-order polynomial nonlinear fitting. The brightness of scatter points from dark to bright means density from low to high.
5.2.2 Results Analysis.
@@ 5.2 Performance Comparison
Fig. 6 shows scatter plots between the predicted scores and the Z-Score MOS of 6 models. We randomly sample 1,000 videos from T2VQA-DB for testing. A better model should have a fitted curve close to the diagonal and have less dispersed scatter points. As shown in Fig. 6, T2VQA also outperforms the others.
5.2.3 Cross-dataset Validation on Sora.
Sora [3] has been considered the SOTA T2V generation model since its release. Since the datasets listed in Tab. 1 haven’t released their subjective scores, we collect the videos generated by Sora To validate the generalization of T2VQA and other models. We collect in total of 48 videos from the official website of Sora. We invite 20 annotators to score the quality of each video. We have all the reference models tested on the Sora videos. T2VQA and other reference VQA models are trained on T2VQA-DB and tested on Sora videos. We report the SROCC, PLCC, KRCC, and RMSE between the model predictions and the ground truth MOSs. The results are listed in Tab. 2.
@@
5.2.4 Qualitative Analysis.
We also conduct a qualitative analysis on three examples with good, fair, and poor quality. We use SimpleVQA [39], BVQA [23], FAST-VQA [47], DOVER [50], and T2VQA to predict their quality. Fig. 7 presents the prompts, video frames, and model predictions on the three examples. Results show that T2VQA has more subjective-aligned predictions. Fig. 6(a) shows an example with a relatively high score. The scene in the video matches the description in the prompt well, and the video frames are clear and consistent. Fig. 6(b) is a medium-level example. Though the video matches the prompt basically, the video frames suffer from blurriness, which is reflected in the MOS and T2VQA’s prediction. Fig. 6(c) shows the worst case among the three examples. The video fails to accurately present the description in the prompt. Besides, it loses consistency between video frames. Although it has a high definition in each frame separately, it still has low scores in both MOS and T2VQA’s prediction.
Refer to caption Prompt: Castle ruins on the hill in the middle of a beautiful landscape. Refer to caption (a)
Refer to caption Prompt: Underwater world with different fishes, corals, and stones. Refer to caption (b)
Refer to caption Prompt: In a hot cast-iron cauldron, the cook pours oil to fry the meat (liver). Refer to caption (c)
Figure 7: Three video examples with good, fair, and poor quality. The deviation between model predictions and MOSs is shown in the histograms.
6 Ablation Studies
@@
6.0.2 Fidelity Encoder.
In T2VQA, we use the Swin-T as the backbone of the fidelity encoder. To investigate the effectiveness of the transformer-based architecture, we conduct the control experiment by using the convolution-based 3D ResNet [13] as the fidelity encoder. The results show that Swin-T has a superior performance to ResNet, validating the effectiveness of the transformer-based architecture.
6.0.3 Fusion and Regression Modules.
@@
Table 3: Performance comparison of ablation studies. The best-performing model is highlighted in each column. [Keys: Bold: the best].
Models Validation
SROCC $\uparrow$ PLCC $\uparrow$ KRCC $\uparrow$ RMSE $\downarrow$
Alignment T2VQA-CLIP 0.7296 0.7347 0.5385 10.5141
Fidelity T2VQA-resnet 0.7610 0.7730 0.5715 9.8152
Fusion T2VQA-cat 0.7734 0.7854 0.5839 9.4034
Regression T2VQA-linear 0.7808 0.7919 0.5891 9.3011
T2VQA-nonlinear 0.7850 0.7983 0.5954 9.1755
Ours T2VQA 0.7965 0.8066 0.6058 9.0221
7 Limitations and Future Work
We investigate the limitations in the proposed T2VQA-DB and T2VQA. For T2VQA-DB, with the fast development of T2V models, models like Sora [3] have achieved stunning results with 1080P, 60s high-fidelity videos. Other current T2V models are not capable of generating such high-fidelity videos. We will expand our dataset with Sora generated videos in future work. For T2VQA, since the datasets listed in Tab. 1 do not release their subjective scores, we will conduct the cross-dataset validation to validate the generalization of T2VQA over other T2V datasets in future work.
8 Conclusion

### v3

@@ 4. Subjective-Aligned Text-to-video Metric
4.1. Text-video Alignment Encoder
@@
Text-video alignment refers to the conformity between the video content and the text description. CLIP (Radford et al., 2021) and BLIP (Li et al., 2022a) are models that have a strong ability for zero-shot text-image matching.  (Wu et al., 2021) first proposes CLIPSim, which uses CLIP to calculate the similarities between text and each frame of the video and then take the average value. However, former works (Chivileva et al., 2023; Huang et al., 2024; Liu et al., 2024, 2023) and results in Tab. 2 show that simply using CLIP or BLIP has a low correlation with the authentic subjective scores. Following works of MLLMs such as BLIP-2 (Li et al., 2023a), InstructBLIP (Dai et al., 2023), and mPLUG-Owl2 (Ye et al., 2023), we use the pre-trained BLIP image encoder as the video frame encoder. We freeze the weights of the image encoder that encodes each frame separately. We use the BLIP text encoder as the alignment encoder. The alignment encoder takes the encoded text and video frame as inputs. They interact through the cross-attention module and the encoder eventually outputs a feature representing the text-frame similarity. We concatenate features from each text-frame pair. Given a set of $N$ video frames $\{v_{i}\in\mathbb{R}^{3\times H\times W}\}_{i=1}^{N}$ and the text prompt $t$, we have:
(2) $f_{b}=cat(\{\mathrm{BLIP}(t,v_{i})\}_{i=1}^{N}).$
4.2. Video Fidelity Encoder
Table 2. Performance of the SOTA models and T2VQA on T2VQA-DB. The best-performing model is highlighted in each column. [Bold: the best].
Type Models T2VQA-DB Validation Sora Testing
SROCC $\uparrow$ PLCC $\uparrow$ KRCC $\uparrow$ RMSE $\downarrow$ SROCC $\uparrow$ PLCC $\uparrow$ KRCC $\uparrow$ RMSE $\downarrow$
zero-shot CLIPSim (Radford et al., 2021) 0.1047 0.1277 0.0702 21.683 0.2116 0.1538 0.1406 18.316
BLIP (Li et al., 2022a) 0.1659 0.1860 0.1112 18.373 0.2126 0.1038 0.1515 18.850
ImageReward (Xu et al., 2024) 0.1875 0.2121 0.1266 18.243 0.0992 0.0415 0.0748 19.494
ViCLIP (Wang et al., 2023b) 0.1162 0.1449 0.0781 21.655 0.2567 0.1844 0.1734 17.982
UMTScore (Liu et al., 2024) 0.0676 0.0721 0.0453 22.559 0.2594 0.0840 0.1680 19.057
finetuned SimpleVQA (Sun et al., 2022) 0.6275 0.6338 0.4466 11.163 0.0340 0.2344 0.0237 16.687
BVQA (Li et al., 2022b) 0.7390 0.7486 0.5487 15.645 0.4235 0.2489 0.2635 17.164
FAST-VQA (Wu et al., 2022) 0.7173 0.7295 0.5303 10.595 0.4301 0.2369 0.2939 17.426
DOVER (Wu et al., 2023e) 0.7609 0.7693 0.5704 9.8072 0.4421 0.2689 0.2757 17.182
Ours T2VQA 0.7965 0.8066 0.6058 9.0221 0.6485 0.3124 0.4874 16.511
@@ 4.2. Video Fidelity Encoder
Video fidelity refers to the perception of distortion from spatial and temporal domains. In the spatial domain, common distortion types include blurriness, noises, low/high contrast, etc. Temporal distortions include jitter, stall, motion blur, etc. In this perspective, the task can be seen as a common VQA task. Swin-T (Liu et al., 2022) has been proven for its excellent ability in various VQA tasks (Wu et al., 2022, 2023e). By using 3D-shifted window-based multi-head self-attention (SW-MSA), Swin-T has a strong ability to analyze videos from spatial and temporal domains. Therefore, we utilize Swin-T as the backbone of the fidelity encoder to extract features that represent video fidelity. Given a video clip $v\in\mathbb{R}^{3\times N\times H\times W}$, we have:
(3) $f_{s}=\mathrm{SWIN}(v).$
4.3. Feature Fusion
@@
Inspired by BLIP-2 (Li et al., 2023a) and InstructBLIP (Dai et al., 2023), after having the features from perspectives of text-video alignment and video fidelity, we design a transformer-based fusion module to fuse those two features. The module includes $N$ blocks with self-attention, cross-attention, and feed-forward layers in each block. The fidelity feature $f_{s}$ first goes through self-attention layers, and then it interacts with the alignment feature $f_{b}$ in cross-attention layers (and every other transformer block). The fusion module helps the model to unify the features from two perspectives to have a more comprehensive understanding of the video characteristics. We initialize the fusion module using $\mathrm{BERT}_{base}$ (Kenton and Toutanova, 2019).
4.4. Quality Regression
@@
LLMs have been proven to be competitive on IQA/VQA tasks (Wu et al., 2024b, 2023f; Zhang et al., 2023c; Wu et al., 2024c). Inspired by them, we also utilize an LLM as the quality regression module in T2VQA. We first design a text instruction prompt, i.e., “Please rate the quality of this video.”, to guide the LLM. The encoded instruction and the fused feature are concatenated as the input of the LLM. Following (Zhang et al., 2023e; Wu et al., 2023f), we supervise the LLM to output one among five ITU-standard (ITU-R, 2002) levels (bad, poor, fair, good, and excellent) to represent the quality of the videos, denoted as <level>. We assign them weights of $1-5$ in order. Since the logit at <level> in LLM is the probability distribution of all tokens, it can be used to represent how the LLM predicts the quality of the video. Therefore, a softmax for each token is calculated and multiplied by its weight. We have the final predicted score as:
(4) $s_{pred}=\sum_{i=1}^{5}i\times softmax(\lambda_{i})=\sum_{i=1}^{5}i\times\frac{e^{\lambda_{i}}}{\sum_{j=1}^{5}e^{\lambda_{j}}},$
@@
5. Experiments and Results
5.1. Implement Details
5.1.1. Train-test splitting.
@@ 5.1. Implement Details
5.1.2. Training Settings.
We utilize the large model of BLIP for the extraction of text-video alignment feature. We initialize the fidelity encoder using Swin-T pre-trained on Kinetics-400 (Kay et al., 2017) dataset. For the LLM, we use the 7B model of Vicuna v1.5 (Zheng et al., 2024), which is fine-tuned from Llama 2 (Touvron et al., 2023). It is worth noting that both BLIP image encoder and LLM are frozen during training.
@@
5.1.3. Loss Function.
(a) BLIP (Li et al., 2022a)
(b) ImageReward (Xu et al., 2024)
(c) ViCLIP (Wang et al., 2023b)
(d) SimpleVQA (Sun et al., 2022)
(e) BVQA (Li et al., 2022b)
(f) FAST-VQA (Wu et al., 2022)
(g) DOVER (Wu et al., 2023e)
(h) T2VQA
Figure 6. Scatter plots of the predicted scores vs. MOSs. The curves are obtained by a four-order polynomial nonlinear fitting. The brightness of scatter points from dark to bright means density from low to high.
Following (Wu et al., 2022), we use differentiable Pearson Linear Correlation Coefficient (PLCC) and rank loss as loss functions. PLCC is a common criterion used for evaluating the correlation between sequences, while rank loss is introduced to help the model distinguish the relative quality of videos better. The final loss function is defined as:
(5) $L=L_{plcc}+\lambda\cdot L_{rank},$
@@
5.1.4. Evaluation Metrics.
Along with PLCC, we include Spearman’s Rank-Order Correlation Coefficient (SROCC), Kendall’s Rank-order Correlation Coefficient (KRCC), and Root Mean Square Error (RMSE) as performance criteria. SROCC, PLCC, and KRCC indicate the prediction monotonicity, while RMSE measures the prediction accuracy. Better models should have larger SROCC, KRCC, and PLCC scores, but conversely for RMSE. Before calculating PLCC, we follow the same procedure in (Group et al., 2000) to map the objective score to the subject score using a four-parameter logistic function.
5.2. Performance Comparison
5.2.1. Reference Algorithms.
We use CLIPSim (Wu et al., 2021), BLIP (Li et al., 2022a), ImageReward (Xu et al., 2024), ViCLIP (Wang et al., 2023b), UMTScore (Liu et al., 2024), SimpleVQA (Sun et al., 2022), BVQA (Li et al., 2022b), FAST-VQA (Wu et al., 2022), and DOVER (Wu et al., 2023e) as the reference algorithms. DOVER is considered the SOTA VQA method to date. CLIPSim averages the similarity values between the text and each video frame. We adopt the same operation on BLIP and ImageReward to tune IQA metrics into VQA metrics. ViCLIP and UMTScore are metrics designed for measuring text-video alignment. SimpleVQA, BVQA, FAST-VQA, and DOVER are models designed for general VQA tasks. We use the pre-trained weights of CLIPSim, BLIP, and ImageReward for zero-shot testing, and we finetune SimpleVQA, BVQA, FAST-VQA, and DOVER on T2VQA-DB. All results are averaged after ten-fold splitting.
5.2.2. Results Analysis.
@@ 5.2. Performance Comparison
Fig. 6 shows scatter plots between the predicted scores and the Z-Score MOSs of BLIP (Li et al., 2022a), ImageReward (Xu et al., 2024), ViCLIP (Wang et al., 2023b), SimpleVQA (Sun et al., 2022), BVQA (Li et al., 2022b), FAST-VQA (Wu et al., 2022), DOVER (Wu et al., 2023e), and T2VQA. We randomly sample 1,000 videos from T2VQA-DB for testing. A better model should have a fitted curve close to the diagonal and have less dispersed scatter points. As shown in Fig. 6, T2VQA also outperforms the others.
5.2.3. Cross-dataset Validation on Sora.
Sora (Brooks et al., 2024) has been considered the SOTA T2V generation model since its release. Since the datasets listed in Tab. 1 haven’t released their subjective scores, we collect the videos generated by Sora To validate the generalization of T2VQA and other models. We collect in total of 48 videos from the official website of Sora. We invite 20 annotators to score the quality of each video. We have all the reference models tested on the Sora videos. T2VQA and other reference VQA models are trained on T2VQA-DB and tested on Sora videos. We report the SROCC, PLCC, KRCC, and RMSE between the model predictions and the ground truth MOSs. The results are listed in Tab. 2.
@@
5.2.4. Qualitative Analysis.
We also conduct a qualitative analysis on three examples with good, fair, and poor quality. We use SimpleVQA (Sun et al., 2022), BVQA (Li et al., 2022b), FAST-VQA (Wu et al., 2022), DOVER (Wu et al., 2023e), and T2VQA to predict their quality. Fig. 7 presents the prompts, video frames, and model predictions on the three examples, and Tab. 3 lists the models’ predictions and MOSs. Results show that T2VQA has more subjective-aligned predictions. Fig. 7a shows an example with a relatively high score. The scene in the video matches the description in the prompt well, and the video frames are clear and consistent. Fig. 7b is a medium-level example. Though the video matches the prompt basically, the video frames suffer from blurriness, which is reflected in the MOS and T2VQA’s prediction. Fig. 7c shows the worst case among the three examples. The video fails to accurately present the description in the prompt. Besides, it loses consistency between video frames. Although it has a high definition in each frame separately, it still has low scores in both MOS and T2VQA’s prediction.
5.3. Ablation Studies
Table 3. T2VQA and other models predictions on 3 example videos. Colored numbers represent the distance between predictions and the ground truths. Red: the best.
Model Prompt 1 Prompt 2 Prompt 3
SimpleVQA (Sun et al., 2022) $72.89_{\color[rgb]{0,1,0}-5.33}$ $61.59_{\color[rgb]{0,1,0}+3.88}$ $49.95_{\color[rgb]{0,1,0}+17.39}$
BVQA (Li et al., 2022b) $73.12_{\color[rgb]{0,1,0}-5.1}$ $60.37_{\color[rgb]{0,1,0}+2.66}$ $45.68_{\color[rgb]{0,1,0}+13.12}$
FAST-VQA (Wu et al., 2022) $85.28_{\color[rgb]{0,1,0}+7.06}$ $44.19_{\color[rgb]{0,1,0}-13.52}$ $38.91_{\color[rgb]{0,1,0}+6.35}$
DOVER (Wu et al., 2023e) $89.11_{\color[rgb]{0,1,0}+10.89}$ $51.93_{\color[rgb]{0,1,0}-5.78}$ $36.66_{\color[rgb]{0,1,0}+4.1}$
T2VQA(Ours) $81.61_{\color[rgb]{1,0,0}+3.39}$ $57.11_{\color[rgb]{1,0,0}-0.6}$ $29.07_{\color[rgb]{1,0,0}-3.49}$
MOS (gt) 78.22 57.71 32.56
(a) Prompt: Castle ruins on the hill in the middle of a beautiful landscape.
(b) Prompt: Underwater world with different fishes, corals, and stones.
(c) Prompt: In a hot cast-iron cauldron, the cook pours oil to fry the meat (liver).
Figure 7. Three video examples with good, fair, and poor quality.
Table 4. Performance comparison of ablation studies. [Keys: Bold: the best].
Models Validation
SROCC $\uparrow$ PLCC $\uparrow$ KRCC $\uparrow$ RMSE $\downarrow$
T2VQA-CLIP 0.7296 0.7347 0.5385 10.5141
T2VQA-resnet 0.7610 0.7730 0.5715 9.8152
T2VQA-cat 0.7734 0.7854 0.5839 9.4034
T2VQA-linear 0.7808 0.7919 0.5891 9.3011
T2VQA-nonlinear 0.7850 0.7983 0.5954 9.1755
T2VQA(Ours) 0.7965 0.8066 0.6058 9.0221
@@ 5.3. Ablation Studies
5.3.2. Fidelity Encoder.
In T2VQA, we use the Swin-T as the backbone of the fidelity encoder. To investigate the effectiveness of the transformer-based architecture, we conduct the control experiment by using the convolution-based 3D ResNet (Hara et al., 2018) as the fidelity encoder. The results show that Swin-T has a superior performance to ResNet, validating the effectiveness of the transformer-based architecture.
5.3.3. Fusion and Regression Modules.
@@
6. Limitations and Future Work
We investigate the limitations in the proposed T2VQA-DB and T2VQA. models like Sora (Brooks et al., 2024) have achieved stunning results with 1080P, 60s high-fidelity videos. Other current T2V models are not capable of generating such high-fidelity videos. We will expand our dataset with Sora generated videos in future work. For T2VQA, since the datasets listed in Tab. 1 do not release their subjective scores, we will conduct the cross-dataset validation to validate the generalization of T2VQA over other T2V datasets in future work.
7. Conclusion